OpenStudy (aonz):

Halp please... Medal to best answer If loga(2) = logb(16) show that b=a^4

OpenStudy (aonz):

@jim_thompson5910 @Kainui Help please

jimthompson5910 (jim_thompson5910):

Hint: use the change of base rule loga(2) = logb(16) log(2)/log(a) = log(16)/log(b)

OpenStudy (aonz):

ahh yes i did that then i couldnt think of a way to go futher :P

jimthompson5910 (jim_thompson5910):

loga(2) = logb(16) log(2)/log(a) = log(16)/log(b) log(2)*log(b) = log(a)*log(16)

jimthompson5910 (jim_thompson5910):

then what

OpenStudy (aonz):

im a slow thinking hold on a min :P

OpenStudy (aonz):

can you even times logs together?

OpenStudy (kainui):

Yeah, there's nothing wrong with multiplying logs together.

OpenStudy (aonz):

log2b =log16a?? Im guessing :P

jimthompson5910 (jim_thompson5910):

no thats not correct

jimthompson5910 (jim_thompson5910):

loga(2) = logb(16) log(2)/log(a) = log(16)/log(b) log(2)*log(b) = log(a)*log(16) log(b)/log(a) = log(16)/log(2) then what?

OpenStudy (aonz):

log(b)*log(2) = log(a)*log(16)

jimthompson5910 (jim_thompson5910):

why would you go back a step?

OpenStudy (aonz):

oh umm my bad

OpenStudy (aonz):

would you multiply everything by log a?

jimthompson5910 (jim_thompson5910):

why not convert use the change of base formula and simplify the left side

OpenStudy (aonz):

loga(b)? =log(16)/log(2)

jimthompson5910 (jim_thompson5910):

now what is log(16)/log(2)

OpenStudy (aonz):

log2(16)?

jimthompson5910 (jim_thompson5910):

use a calculator if you're not sure notice that log(16)/log(2) is the same as log2(16) yes

jimthompson5910 (jim_thompson5910):

and that is?

OpenStudy (aonz):

4

OpenStudy (aonz):

AHH I GET IT NOW :P :D

jimthompson5910 (jim_thompson5910):

so you would get loga (b) = 4

jimthompson5910 (jim_thompson5910):

ok good